OpenStudy (anonymous):

What is the numerator of the fraction equivalent to 25/135 that has a denominator of 27? 15 10 5 3

OpenStudy (anonymous):

\[\frac{ 25/5 }{ 135/5 } \]\[\frac{ 5 }{ 27 }\]

OpenStudy (michele_laino):

being \(27 \cdot 5=135\) you have to divide numerator by \(5\)